OpenStudy (anonymous):

PLEASE HELP f^1(3) when f(x) = (2x+3)/5

OpenStudy (jhannybean):

Did you mean \(f^{-1}(3)\)?

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

First you want to set f(x) = y Can you do that for me?

OpenStudy (anonymous):

Y = f(x) (2x+3)/5 ?????

OpenStudy (jhannybean):

I meant replace the f(x) with y =...

OpenStudy (anonymous):

Oh okay then it owuld be y = (2x+3)/5

OpenStudy (jhannybean):

Awesome :) Now switch `y` with `x` , and rewrite the equation.

OpenStudy (anonymous):

x = (2y+3)/5

OpenStudy (jhannybean):

Great :) We want to solve for `y`, so the first thing we would need to do is multiply both sides of the equation by `5`. Can you show me what the equation becomes when you do that?

OpenStudy (anonymous):

5x = (2y+3)

OpenStudy (jhannybean):

Now we want to subtract -3 from both sides. What does your equation become?

OpenStudy (anonymous):

5x-3 = 2y

OpenStudy (jhannybean):

And now divide both sides of your equation by 2.

OpenStudy (anonymous):

5/2x-3/2 = y

OpenStudy (jhannybean):

And now we have found our inverse function :)\[f^{-1}(x) = \frac{5x-3}{2} \iff \frac{5x}{2} -\frac{3}{2}\]

OpenStudy (anonymous):

Oh that makes so much sense, so is that it? Or are there more steps?

OpenStudy (jhannybean):

Now we want to find \(f^{-1}(3)\) , so all you have to do now is plug in `3` for x and solve :)

OpenStudy (jhannybean):

What do you get?

OpenStudy (anonymous):

15/2 - 3/2 so that would be 12/2 so then it would be 6???

OpenStudy (jhannybean):

Yay! I got the same :D

OpenStudy (jhannybean):

Good job ^^

OpenStudy (anonymous):

Okay awesome, so just to clarify would it be x = 6 or y = 6???

OpenStudy (jhannybean):

I believe it is x.

OpenStudy (jhannybean):

One thing I forgot to mention is we started out with \[f(x) = \frac{2x+3}{5}\]We rewrote it as \[y=\frac{2x+3}{5}\]Swapped variables.\[x=\frac{2y+3}{5}\]solved for y again.\[y=\frac{5x-3}{2}\]and then rewrote it once again as \(f^{-1}(x)\)\[f^{-1}(x) = \frac{5x-3}{2}\]To check which variable we are solving for, we could take the inverse of the original function, this will leave us with the function itself, either in x, or y. To check: \[f^{-1}(f(x)) = x\]